OpenStudy (anonymous):

what is the solution to the rational equation (5/(x^2+3x+2)) + (1/(x+2)) = (1/(3x+3))

OpenStudy (jiteshmeghwal9):

\[\frac{5}{x^2+3x+2}=\frac{1}{3x+3}-\frac{1}{x+2}\]

OpenStudy (jiteshmeghwal9):

now factor L.H.S. & then cross multiply to gt ur answer;)

OpenStudy (anonymous):

but i need to know what x equals